Calculate the [OH–] in a 1.0 × 10–3 M solution of the weak base methylamine

(CH3NH2), whose Kb = 4.4 × 10–4. What percentage of the base has dissociated?


1
Expert's answer
2021-03-05T03:28:05-0500

CH3NH2 + H2O = CH3NH3+ + OH-

Kb = ([CH3NH3+][OH-])/[CH3NH2] = [OH-]2/[CH3NH2] = x2/(1*10-3 - x) = 4.4*10-4

Solving the equation we get [OH-] = x = 0.000478856 M and percent dissociation x*100% = 0.0479%